a force of magnitude 200 n is directed from a(0 3 0) to b(0 0 4) the z component of force is


1
Expert's answer
2021-08-03T11:41:29-0400

Explanations & Calculations




  • If you look carefully you can figure out that even though the coordinates define a 3d vector space, the given situation is restrained to the y-z plane, which makes the process much easier.
  • Then the angle alpha can be calculated & then the force can be resolved along the needed direction.

"\qquad\qquad
\begin{aligned}
\small \tan \alpha &=\small \frac{4}{3}\\\\

\small \therefore \,F_z&=\small 200N\times \sin \alpha=200\times \frac{4}{5}\\
&=\small \bold{160\,N} 
\end{aligned}"